Find the domain of each relation if the range is {0, 4, 36} . There are two equations 1. y=x squared 2. y= |2x - 4|
Answer by stanbon(75887)

Find the domain of each relation if the range is {0, 4, 36} .
There are two equations
1. y=x squared
Solve each of the following equation for x to determine the domain.
x^2 = 0
x^2 = 4
X^2 = 36
---
Ans: Domain = {0,+-2,+-6}
-------------------------------
2. y= |2x - 4|
|2x-4| = 0
Solve 2x-4 = 0
x = 2
===================
|2x-4| = 4
Solve 2x-4 = 4 or 2x-4 = -4
x = 4 or x = 0
===================
|2x-4| = 36
Solve 2x-4 = 36 or 2x-4 = -36
x = 20 or x = -16
--
Ans: Domain = {0,2,4,20,-16}
===============================
Cheers,
Stan H.
---